# Painless Activation Steering: An Automated, Lightweight Approach for Post-Training Large Language Models

### v2

@@ Abstract
Language models (LMs) are typically post-trained for desired capabilities and behaviors via weight-based or prompt-based steering, but the former is time-consuming and expensive, and the latter is not precisely controllable and often requires manual trial-and-error. While activation steering (AS) promises a cheap, fast, and controllable alternative to the two existing post-training methods, current AS techniques require hand-crafted prompt pairs or labor-intensive feature annotation, making them more inconvenient than the plug-and-play methods such as Reinforcement Learning (RL) and Supervised Fine-Tuning (SFT). We introduce Painless Activation Steering (PAS), a family of fully automated methods that make AS readily usable with any given labeled dataset, with no need for prompt construction, feature labeling, or human intervention. We evaluate PAS on three open-weight models (Llama3.1-8B-Instruct, DeepSeek-R1-Distill-8B, and Nous-Hermes-2) and 18 tasks; we find that PAS reliably improves performance for behavior tasks, but not for intelligence-oriented tasks. The introspective variant (iPAS) delivers the strongest causal steering effects (10.1% on Bias, 5.2% on Morality, and 34.8% on Alignment). We also show PAS delivers additional gains on top of In-Context Learning (ICL) and SFT. PAS constructs a fast, lightweight activation vector that can be cheaply trained, easily stored, and activated at will. Our results provide a characterization of where AS helps, where it fails, and how to deploy it as a practical, automated LM post-training option.
Figure 1: Average causal steering effects on behavior tasks. Each bar reports the mean improvement in test accuracy relative to the unsteered baseline, averaged across 3 models and 15 trials. Colored bars correspond to different steering methods. Black vertical lines denote 95% confidence intervals.
Figure 2: iPASwo (introspective PAS-wrong only) pipeline; prompts are built from the model’s own errors. (1) Run the raw LM on the training split and partition items into correct vs. incorrect. (2) From the incorrect items, build positive prompts using the ground-truth answers and negative prompts using the model’s chosen (incorrect) answers. (3) Compute a steering vector $a^{*}$ as the mean activation difference between the two prompt sets at a chosen layer $\ell$ and target steer_targ. (4) At inference, inject this vector (with strength $\lambda$) to obtain the activation-steered LM. (5) Evaluate the steered model on the held-out test split.
@@ 1 Introduction
To modify the behaviors of pre-trained Language Models (LMs), one typically either changes the weights, such as Reinforcement Learning (RL) (Ouyang et al., 2022) and Supervised Fine-Tuning (SFT) (Radford et al., 2018), or the prompts, such as In-Context Learning (ICL) and Context Engineering (Brown et al., 2020; Zhao et al., 2021; Agrawal et al., 2025). Recent studies in mechanistic interpretability and representation engineering have shown that model behaviors can be modified by interventions on the activations (Turner et al., 2023; Panickssery et al., 2024; Zou et al., 2025; Meng et al., 2022). During inference, AS injects steering vectors into the internal neuron activations without changing the weights or prompts. Recent works on AS suggest that inference-time causal Activation Steering (AS) potentially offers a cheap and flexible third avenue for model post-training.
Problem: Current Activation Steering Methods Are Human-Dependent and Impractical
@@
While there have been sporadic examples showcasing the applicability of AS to customized settings (e.g., in eliciting or shortening Chains of Thoughts (Zhang & Viteri, 2025; Azizi et al., 2025)), there is neither an automated way of cheaply applying AS to arbitrary tasks nor a clear understanding of the contexts in which AS is suitable. Prior works in AS rely on either pre-labeled Sparse Autoencoders (SAE) (Smith et al., 2025; Soo et al., 2025; Bayat et al., 2025) or manually constructed pairs of prompts with positive and negative examples of the desired behaviors, as is the case for traditional Activation Steering methods (Panickssery et al., 2024; Chen et al., 2025; Lee et al., 2025; Turner et al., 2023) and probe-based activation construction methods (O’Neill et al., 2025; Goldowsky-Dill et al., 2025). Humans or frontier LMs are needed to identify what each sparse feature represents (in the former case) or to build and filter high-quality prompt pairs representative of each desired behavior (in the latter case). These carefully collected SAE features and crafted prompts are costly to build and useful only in limited scenarios, thus restricting the practical applicability of traditional AS methods. This raises the following question: Do there exist AS methods that are both human-independent and adaptive to arbitrary models and arbitrary tasks?
Contribution: An Automated Pipeline for Activation Steering
@@
Our answer is yes. In this paper, we present Painless AS (PAS), a family of automation-friendly, human-independent methods that overcome the scalability limitations of existing AS approaches. PAS removes humans and external LM calls from the loop, allowing it to run on any labeled dataset, just like SFT or RL.
Finding: PAS Effectively Changes Behaviors, Fails to Improve Intelligence, and Rarely Causes Catastrophic Forgetting
@@
We evaluate PAS across a variety of tasks to characterize its strengths and limitations. We discover that PAS is effective for behavior tasks including changing Bias and Sentiments (10 tasks), Moral Preferences (3 tasks), and Alignment (2 tasks), but PAS provides little benefit on intelligence (3 tasks) where the model’s knowledge and reasoning ability are evaluated. PAS usually does not lead to catastrophic forgetting in our experiments—we find that it decreases general capabilities by less than 2% on 13 out of 15 behavior tasks. Among the 18 diverse tasks, the proposed method boosts accuracy by 10.1% on Bias, 5.2% on Morality, 34.8% on Alignment, but yields no significant gain on Intelligence tasks.
Finding: Introspection Performs Best
@@
We find that introspective PAS (iPAS) variants perform best. In iPAS, each LM identifies weaknesses in the training split, internalizes lessons from its past mistakes, and then applies them to novel problems—analogous to the student who prepares for a test by reviewing errors from homework. iPAS extends the error-driven, contrastive, and self-supervised learning traditions in reasoning (Zelikman et al., 2022), vision (Chen et al., 2020), and agentic scaffolding (Shinn et al., 2023). In doing so, it overcomes the rigidity of static prompt pairs (Chen et al., 2025), which cannot adapt steering vectors to a model’s unique weaknesses.
Finding: PAS Complements Weight- and Prompt-Based Steering
@@
We study how PAS can compete with or complement existing weight-based and prompt-based post-training methods. Compared to weight-based steering, PAS enjoys three advantages.
@@
Compared to prompt-based steering, PAS enjoys two advantages.
@@
    Higher Improvement Ceiling: PAS complements prompt-based steering. We find that PAS complements ICL by yielding additional gains, with average improvements of 2.9% on Bias tasks, 3.4% on Morality tasks, and 17.0% on Alignment tasks beyond the ICL baseline.
  2. 2.
@@
Practical Takeaway
For customization or personalization tasks, PAS is an effective lightweight option. We hope that researchers and practitioners consider using PAS to complement weight- and prompt-based post-training methods.
@@ 2.1 Notation and Preliminaries for Post-Training & Activation Steering
Evaluation Dimensions and Post-Training
Suppose we have $K$ distinct evaluation dimensions of interest, with measured task performance scores $Y_{k}(M)\in\mathbb{R}^{+}$ for $k\in[K]$. To post-train a model, we begin by specifying a partition $(\{k\},\Phi,\Psi)$ of the set $[K]$, where $k$ is the target evaluation dimension, $\Phi$ is the set of control evaluation dimensions, and $\Psi$ is the set of unconstrained evaluation dimensions. For each partition $(\{k\},\Phi,\Psi)$, a post-training method specifies a pipeline that takes an LM ($M$) and creates a modified LM ($\overline{M}$) satisfying
$\forall\phi\in\Phi,\,\,\mathbb{E}\left[Y_{\phi}(\overline{M})-Y_{\phi}(M)\right]\geq-\epsilon_{\phi}\quad\text{and}\quad\mathbb{E}\left[Y_{k}(\overline{M})-Y_{k}(M)\right]\geq\epsilon_{k}.$
@@
Steer Target and Injection of Steering Vectors
As a lightweight post-training method, AS builds $\overline{M}$ by injecting at inference time the steering vector $a^{*}$ to layer $\ell$ of the raw model $M$ with steering strength $\lambda$. For every token produced, the activation at the steer_targ of layer $\ell$ is modified by
$a_{\ell}(\texttt{steer\_targ})\leftarrow a_{\ell}(\texttt{steer\_targ})+\lambda\cdot a^{*}.$ (1)
Here, steer_targ denotes the location in the transformer architecture where the hook is attached and activations are collected and injected. The modifiable parameters of AS are the activation vector $a^{*}$, the steering layer $\ell$, the steering strength $\lambda$, and the target location steer_targ.
Construction of Steering Vectors
We begin the construction of the steering vector $a^{*}$ by recording activations of the raw model $M$ at the last token of a variety of prompts. Since the model is autoregressive, the activation at the last token reflects the cumulative effect of the entire prompt. [^2] This is the standard practice in the literature. We conducted some preliminary experiments on alternative ways of activation pooling or weighing (e.g., computing the averaged activations over all tokens of the prompt). The results were unpromising and we abandoned this line of inquiry. For each evaluation dimension $k$, we construct two sets of prompts, $\mathbf{P}^{+}_{k}:=\left\{p_{1}^{+}(k),\dots,p_{n^{+}_{k}}^{+}(k)\right\}$ and $\mathbf{P}^{-}_{k}:=\left\{p_{1}^{-}(k),\dots,p_{n^{-}_{k}}^{-}(k)\right\}$. The positive prompt set contains $n^{+}_{k}$ examples of the desired behavior, while the negative prompt set contains $n^{-}_{k}$ examples of the opposite behavior. The activation vector is constructed as the mean activation difference of the model at the corresponding layer and steer target location when supplied these prompts, i.e.,
$a^{*}_{\ell}(\texttt{steer\_targ},k):=\sum_{j=1}^{n_{k}^{+}}\frac{a_{\ell}(\texttt{steer\_targ};p_{j}^{+}(k))}{n_{k}^{+}}-\sum_{j=1}^{n_{k}^{-}}\frac{a_{\ell}(\texttt{steer\_targ};p_{j}^{-}(k))}{n_{k}^{-}}.$ (2)
@@ 2.2 Painless Activation Steering: Automated Strategies for Steering Vectors Extraction
How is $(\mathbf{P}^{+}_{k},\mathbf{P}^{-}_{k})$ built? Unlike existing methods (Panickssery et al., 2024; Chen et al., 2025), which require the careful construction of $(\mathbf{P}^{+}_{k},\mathbf{P}^{-}_{k})$’s for each model and target dimension, PAS constructs the prompt sets automatically. We achieve this by first running the raw model $M$ on the training split of a dataset (Fig. 2). We then use the correctly and incorrectly solved tasks to construct $(\mathbf{P}^{+}_{k},\mathbf{P}^{-}_{k})$. We also automatically search for the optimal intervention layer $\ell$ and the optimal steering strength $\lambda$ on a validation split. The resultant model $\overline{M}$ is then evaluated on the test split.
There are several ways to construct $(\mathbf{P}^{+}_{k},\mathbf{P}^{-}_{k})$ from the performance of the model on the training split (Table 1). The PAS (Full MCQ) variant uses full multiple-choice questions—those the model answered correctly form the positive prompts, and those it answered incorrectly form the negative prompts. The introspective variants (iPAS) tailor the prompts to each model’s specific weaknesses. We adaptively construct the positive and negative prompt sets so that the steering vector $a^{*}$ directs the model away from the specific mistakes it made in the training set.
@@
Positive Prompt Set: What is the color of a tiger’s fur? A: Blue. B: Red. C: Orange.
Negative Prompt Set: What is the capital of France? A: Paris. B: London. C: Rome.
Method 2 (iPAS-All)
@@
Positive Prompt Set: What is the color of a tiger’s fur? Orange.
Negative Prompt Set: What is the capital of France? London.
Method 3 (iPAS-Wrong-Only)
@@
Positive Prompt Set: What is the capital of France? Paris.
Negative Prompt Set: What is the capital of France? London.
Strategy Name Positive Prompt Set $\mathbf{P}^{+}_{k}$ Negative Prompt Set $\mathbf{P}^{-}_{k}$
@@ 3.1 Experiment Design
We compare the unsteered baseline with three steering methods, iPAS–all (iPASa), iPAS–wrong‑only (iPASwo), and PAS–full‑MCQ (PASf), under various experimental settings. For each evaluation dimension $k$ and each method, we define the causal steering effect as the expected difference in performance on unseen test data relative to the raw model,
@@
where $\overline{M}\in\{\text{iPASa}(M,k),\text{iPASwo}(M,k),\text{PASf}(M,k)\}$, that is, the model steered by the corresponding PAS strategy for target evaluation dimension $k$. Here, $Y_{k}(\overline{M})$ denotes the performance of the steered model $\overline{M}$ and $Y_{k}(\mathrm{M})$ denotes the performance of the raw model, both on evaluation dimension $k$. We systematically use greedy decoding to extract model answers, and the source of randomness is the choice of the random seeds. For each experimental setting, we run 15 trials with different random seeds per task and conduct a paired $t$-test [^3] We used one-sided tests as our hypotheses are directional $(\text{steering}\geq\text{baseline})$ to test the null hypothesis $H_{0}:\Delta_{\overline{M}}=0$.
3.2 Causal Steering Effect on behavior & Intelligence-Oriented Tasks

### v3

@@ Abstract
Language models (LMs) are typically post-trained for desired capabilities and behaviors via weight-based or prompt-based steering, but weight updates are expensive and prompt-based methods can be brittle and hard to control. Activation steering (AS) is a cheap, fast, and controllable alternative, yet prior methods often rely on hand-crafted prompt pairs or labor-intensive feature annotation. We introduce Painless Activation Steering (PAS), a fully automated family of methods that turns any labeled dataset into steering vectors with no prompt construction, feature labeling, or human intervention. Across three open-weight models (Llama3.1-8B-Instruct, DeepSeek-R1-Distill-8B, and Nous-Hermes-2) and 18 tasks, PAS reliably improves behavior-task performance but not intelligence-oriented tasks. The introspective variant (iPAS) delivers the strongest steering effects (10.1% on Bias, 5.2% on Morality, and 34.8% on Alignment) and provides additional gains on top of In-Context Learning (ICL) and Supervised Fine-Tuning (SFT). PAS produces lightweight activation vectors that are cheap to train, easy to store, and can be activated on demand, clarifying where AS helps and where it fails.
Painless Activation Steering:
An Automated, Lightweight Approach for Post-Training Large Language Models
Sasha Cui Department of Statistics and Data Science Yale University New Haven, CT, USA sasha.cui@yale.edu       Zhongren Chen Department of Statistics and Data Science Yale University New Haven, CT, USA zhongren.chen@yale.edu
@@ 1 Introduction
To modify the behaviors of pre-trained Language Models (LMs), one typically either changes the weights, such as Reinforcement Learning (RL) (Ouyang et al., 2022) and Supervised Fine-Tuning (SFT) (Radford et al., 2018), or the prompts, such as In-Context Learning (ICL) and Context Engineering (Brown et al., 2020; Zhao et al., 2021; Agrawal et al., 2025). Recent studies in mechanistic interpretability and representation engineering have shown that model behaviors can be modified by interventions on the activations (Turner et al., 2023; Panickssery et al., 2024; Zou et al., 2025; Meng et al., 2022). During inference, Activation Steering (AS) injects steering vectors into the internal neuron activations without changing the weights or prompts. Recent works suggest that inference-time AS potentially offers a cheap and flexible third avenue for model post-training.
Problem: Current Activation Steering Methods Are Human-Dependent and Impractical There have been sporadic examples showcasing the applicability of AS to customized settings (e.g., in eliciting or shortening Chains of Thoughts (Zhang and Viteri, 2025; Azizi et al., 2025)).
@@
However, there is neither an automated way of cheaply applying AS to arbitrary tasks nor a clear understanding of the contexts in which AS is suitable. Prior works in AS rely on either pre-labeled Sparse Autoencoders (SAE) (Smith et al., 2025; Soo et al., 2025; Bayat et al., 2025) or manually constructed pairs of prompts with positive and negative examples of the desired behaviors, as is the case for traditional Activation Steering methods (Panickssery et al., 2024; Chen et al., 2025; Lee et al., 2025; Turner et al., 2023) and probe-based activation construction methods (O’Neill et al., 2025; Goldowsky-Dill et al., 2025). Humans or frontier LMs are needed to identify what each sparse feature represents (in the former case) or to build and filter high-quality prompt pairs representative of each desired behavior (in the latter case). These carefully collected SAE features and crafted prompts are costly to build and useful only in limited scenarios, thus restricting the practical applicability of traditional AS methods. This raises the following question: Do there exist AS methods that are both human-independent and adaptive to arbitrary models and a broad range of labeled tasks?
Figure 1: iPASwo (introspective PAS-wrong only) pipeline; prompts are built from the model’s own errors. (1) Run the raw LM on the training split and partition items into correct vs. incorrect. (2) From the incorrect items, build positive prompts using the ground-truth answers and negative prompts using the model’s chosen (incorrect) answers. (3) Compute a steering vector $a^{*}$ as the mean activation difference between the two prompt sets at a chosen layer $\ell$ and target st. (4) At inference, inject this vector (with strength $\lambda$) to obtain the activation-steered LM. (5) Evaluate the steered model on the held-out test split.
@@
Contribution: An Automated Pipeline for Activation Steering Our answer is yes. In this paper, we present Painless AS (PAS), a family of automation-friendly, human-independent methods that overcome the scalability limitations of existing AS approaches. PAS removes humans and external LM calls from the loop, allowing it to run on any labeled dataset, just like SFT or RL.
@@
Finding: PAS Effectively Changes Behaviors, Fails to Improve Intelligence, and Rarely Causes Catastrophic Forgetting We evaluate PAS across a variety of tasks to characterize its strengths and limitations. We discover that PAS is effective for behavior tasks including changing Bias and Sentiments (10 tasks), Moral Preferences (3 tasks), and Alignment (2 tasks), but PAS provides little benefit on intelligence (3 tasks) where the model’s knowledge and reasoning ability are evaluated. PAS usually does not lead to catastrophic forgetting in our experiments. Among the 18 diverse tasks, the proposed method boosts accuracy by 10.1% on Bias, 5.2% on Morality, 34.8% on Alignment, but yields no significant gain on Intelligence tasks.
@@
Finding: Introspection Performs Best We find that introspective PAS (iPAS) variants perform best. In iPAS, each LM identifies weaknesses in the training split, internalizes lessons from its past mistakes, and then applies them to novel problems—analogous to the student who prepares for a test by reviewing errors from homework. iPAS extends the error-driven, contrastive, and self-supervised learning traditions in reasoning (Zelikman et al., 2022), vision (Chen et al., 2020), and agentic scaffolding (Shinn et al., 2023). In doing so, it overcomes the rigidity of static prompt pairs (Chen et al., 2025), which cannot adapt steering vectors to a model’s unique weaknesses.
@@
Finding: PAS Complements Weight- and Prompt-Based Steering We study how PAS can compete with or complement existing weight-based and prompt-based post-training methods. PAS is favourable to weight-based steering in three ways.
@@
PAS is favourable to prompt-based steering in two ways.
@@
We hope that researchers and practitioners consider using PAS to complement weight- and prompt-based post-training methods.
@@ 2 Methods
Strategy Name Pos Prompt Set $\mathbf{P}^{+}_{k}$ Neg Prompt Set $\mathbf{P}^{-}_{k}$
@@ 2.1 Notation and Preliminaries for Post-Training & Activation Steering
Evaluation Dimensions and Post-Training Suppose we have $K$ distinct evaluation dimensions of interest, with measured task performance scores $Y_{k}(M)\in\mathbb{R}^{+}$ for $k\in[K]$. To post-train a model, we begin with a partition $(\{k\},\Phi,\Psi)$ of the set $[K]$, where $k$ is the target evaluation dimension, $\Psi$ is the set of unconstrained evaluation dimensions, for which performance decreases are irrelevant to the use case. For each partition $(\{k\},\Phi,\Psi)$, a post-training method specifies a pipeline turning an LM, $M$, into a modified LM, $\overline{M}$, satisfying
$\forall\phi\in\Phi,\,\,\mathbb{E}\left[Y_{\phi}(\overline{M})-Y_{\phi}(M)\right]$ $\geq-\epsilon_{\phi}$
$\text{and}\quad\mathbb{E}\left[Y_{k}(\overline{M})-Y_{k}(M)\right]$ $\geq\epsilon_{k}.$
@@
Steer Target and Injection of Steering Vectors As a lightweight post-training method, AS builds $\overline{M}$ by injecting at inference time the steering vector $a^{*}$ to layer $\ell$ of the raw model $M$ with steering strength $\lambda$. For every token produced, the activation at the st of layer $\ell$ is modified by
$a_{\ell}(\texttt{st})\leftarrow a_{\ell}(\texttt{st})+\lambda\cdot a^{*}.$ (1)
Here, st denotes the location in the transformer architecture where the hook is attached and activations are collected and injected. The modifiable parameters of AS are the activation vector $a^{*}$, the steering layer $\ell$, the steering strength $\lambda$, and the target location st.
Construction of Steering Vectors We begin the construction of the steering vector $a^{*}$ by recording activations of the raw model $M$ at the last token of a variety of prompts. Since the model is autoregressive, the activation at the last token reflects the cumulative effect of the entire prompt. [^2] This is the standard practice in the literature. We conducted some preliminary experiments on alternative ways of activation pooling or weighing (e.g., computing the averaged activations over all tokens of the prompt). The results were unpromising and we abandoned this line of inquiry. For each evaluation dimension $k$, we construct two sets of prompts, $\mathbf{P}^{+}_{k}:=\left\{p_{1}^{+}(k),\dots,p_{n^{+}_{k}}^{+}(k)\right\}$ and $\mathbf{P}^{-}_{k}:=\left\{p_{1}^{-}(k),\dots,p_{n^{-}_{k}}^{-}(k)\right\}$. The Pos Prompt Set contains $n^{+}_{k}$ examples of the desired behavior, while the Neg Prompt Set contains $n^{-}_{k}$ examples of the opposite behavior. The activation vector is constructed as the mean activation difference of the model at the corresponding layer and steer target location when supplied these prompts, i.e.,
$a^{*}_{\ell}(\texttt{st},k)$ $:=\sum_{j=1}^{n_{k}^{+}}\frac{a_{\ell}(\texttt{st};p_{j}^{+}(k))}{n_{k}^{+}}-\sum_{j=1}^{n_{k}^{-}}\frac{a_{\ell}(\texttt{st};p_{j}^{-}(k))}{n_{k}^{-}}.$ (2)
@@ 2.2 Painless Activation Steering: Automated Strategies for Steering Vectors Extraction
How is $(\mathbf{P}^{+}_{k},\mathbf{P}^{-}_{k})$ built? Unlike existing methods (Panickssery et al., 2024; Chen et al., 2025), which require the careful construction of $(\mathbf{P}^{+}_{k},\mathbf{P}^{-}_{k})$’s for each model and target dimension, PAS constructs the prompt sets automatically. We achieve this by first running the raw model $M$ on the training split of a dataset (Figure 1). We then use the correctly and incorrectly solved tasks to construct $(\mathbf{P}^{+}_{k},\mathbf{P}^{-}_{k})$. We also automatically search for the optimal intervention layer $\ell$ and the optimal steering strength $\lambda$ on a validation split. The resultant model $\overline{M}$ is then evaluated on the test split.
There are several ways to construct $(\mathbf{P}^{+}_{k},\mathbf{P}^{-}_{k})$ from the performance of the model on the training split (Table 1). The PAS (Full MCQ) variant uses full multiple-choice questions—those the model answered correctly form the positive prompts, and those it answered incorrectly form the negative prompts. The introspective variants (iPAS) tailor the prompts to each model’s specific weaknesses. We adaptively construct the positive and Neg Prompt Sets so that the steering vector $a^{*}$ directs the model away from the specific mistakes it made in the training set.
@@
Pos Prompt Set: What is the color of a tiger’s fur? A: Blue. B: Red. C: Orange.
Neg Prompt Set: What is the capital of France? A: Paris. B: London. C: Rome.
Method 2.2 (iPAS-All).
@@
Pos Prompt Set: What is the color of a tiger’s fur? Orange.
Neg Prompt Set: What is the capital of France? London.
Method 2.3 (iPAS-Wrong-Only).
@@
Pos Prompt Set: What is the capital of France? Paris.
Neg Prompt Set: What is the capital of France? London.
Figure 2: PAS consistently improves behavior-task performance. Bars show mean test-accuracy improvements over the unsteered baseline, averaged across three models and 15 trials. Error bars denote 95% confidence intervals. Across Bias, Morality, and Alignment tasks, PAS variants produce broadly positive gains and generally outperform Contrastive Activation Addition (CAA) (Panickssery et al., 2024).
@@ 3 Experiment Design & Results
We organize the experiments around four questions: whether PAS improves target behavior tasks, whether these gains extend to intelligence-oriented tasks, whether PAS remains robust across steering targets and hyperparameters, and whether it complements ICL/SFT without substantial catastrophic forgetting.
@@ 3.1 Experiment Design
We compare the unsteered baseline with three steering methods, iPAS–all (iPASa), iPAS–wrong-only (iPASwo), and PAS–full-MCQ (PASf), under various experimental settings. For each evaluation dimension $k$ and each method, we define the steering effect as the expected difference in performance on unseen test data relative to the raw model,
@@
where $\overline{M}\in\{\text{iPASa}(M,k),\text{iPASwo}(M,k),\text{PASf}(M,k)\}$, that is, the model steered by the corresponding PAS strategy for target evaluation dimension $k$. Here, $Y_{k}(\overline{M})$ denotes the performance of the steered model $\overline{M}$ and $Y_{k}(\mathrm{M})$ denotes the performance of the raw model, both on evaluation dimension $k$. We systematically use greedy decoding to extract model answers, and the source of randomness is the choice of the random seeds. For each experimental setting, we run 15 random trials per task and conduct a paired $t$-test [^3] We used one-sided tests as our hypotheses are directional $(\text{steering}\geq\text{baseline})$ to test the null hypothesis $H_{0}:\Delta_{\overline{M}}=0$. Unless specified otherwise, we always use the ratio $n_{\text{train}}:n_{\text{val}}:n_{\text{test}}=3:1:1$ with $n_{\text{train}}\leq 2400$ cutoff. The unreported dimensions are considered “unconstrained” and belong to $\Phi$.
3.2 Steering Effect on Behavior & Intelligence Tasks